OpenStudy (anonymous):

can some one help me diferentiate x^1/2+x^-4/3

OpenStudy (anonymous):

in respect to what?

OpenStudy (anonymous):

to y

OpenStudy (anonymous):

\[y = x ^{1/2} + x ^{-4/3} \] is this the original function?

OpenStudy (anonymous):

no the original function is y=sqrt of x + 1/cube root of x^4

OpenStudy (anonymous):

write it as an equation because now im conufse

OpenStudy (anonymous):

\[\sqrt{x}+1/\sqrt[3]{x^4}\]

OpenStudy (anonymous):

dy/dx = 1/2 x^(-1/2) -4/3 x^-7/3

OpenStudy (anonymous):

rewrite it as \[y= x ^{1/2} + 1 /(x ^{4/3})\] \[y= x ^{1/2} + (x ^{-4/3})\] \[dy/dx= (1/2)x ^{-1/2} + (-4/3)x ^{-7/3}\]

OpenStudy (anonymous):

understand how to get it?

OpenStudy (anonymous):

thats the first answer i got but i dont uderstand how i got it

OpenStudy (anonymous):

i understand the first step but am lost at the second

OpenStudy (anonymous):

are we using both the product rule and quotient rule here?

OpenStudy (anonymous):

no, just differentiate each term separately

OpenStudy (anonymous):

its the chain rule for example when you find the derivative of \[x^a\] the answer is \[a * x ^{a-1}\]

OpenStudy (anonymous):

@Romero, I dont think its the chain rule since this is not a composite function this question only requires you to do basic differentiation using the rule "y=(a)x^n" differentiates to "dy/dx= (a)(n)x^(n-1)

OpenStudy (anonymous):

sorry for not getting the word right but the concept is the same in finding the derivative.

OpenStudy (anonymous):

see i actually got what romero got but tried to the product rule after

OpenStudy (anonymous):

You dont need to use the product or quotient rule here. That's only when you have 2 variables multiplying to dividing.

OpenStudy (anonymous):

yeah i just saw that

OpenStudy (anonymous):

thank you so much man i really appreciate it trying to cram for exam am going to have in a couple of days and this really helps

OpenStudy (anonymous):

Sure thing man. I think the problem you're having is rewriting out square roots. Here this might help.\[\sqrt[a]{x^b} = x ^{(b/a)}\] if you have no number for a then a =2 also \[1/x^a = 1*x ^{-a}\] It was hard for me to see it at first but in time this will become second hand to you.

OpenStudy (anonymous):

do you think with couple of more problems i think i got how to do them but i dont know if they are right

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

y=cos(tanx) have to deferentiate

OpenStudy (anonymous):

you use the chain rule so you should get y'= sin(tan x) * sec^2 x

OpenStudy (anonymous):

why do i use the product rule here

OpenStudy (anonymous):

no you dont use the product rule you only have one x within tan that happens to be inside cos

OpenStudy (anonymous):

actually its suppose to be y'= -sin(tan x) * sec^2 x

OpenStudy (anonymous):

With this problem you need to use the chain rule because its a composite function. The chain rule is, if y=f(u) \[dy/dx = dy/du*du/dx\] Let u=tanx y=cos(u) differentiates to -sin(u) tanx differentiates to \[\sec ^{2}x\] therefore dy/dx= -sin(tanx)*\[\sec ^{2}x\]

OpenStudy (anonymous):

when y= cos(a) y' = cos(a)' * a' in this case a = tan(x) so y'= - sin(tan(x)) * tan(x)' where tan(x)'= sec^2 (x) so y'= - sin (tanx) * sec^2 x

OpenStudy (anonymous):

oh ok i am getting it

OpenStudy (anonymous):

how about 3x-2/\[\sqrt{2x+1}\]

OpenStudy (anonymous):

Use the quotient rule.

OpenStudy (anonymous):

how do i diferentiate \[\sqrt{2x+1}\]

OpenStudy (anonymous):

y= a/b where y'= (a' * b - b' * a)/ b^2 in this case a = 3x -2 \[b= \sqrt{2x + 1}\]

OpenStudy (anonymous):

so a'=3 right

OpenStudy (anonymous):

yes a'=3, b'=1/2(2x+1)^(-1/2)

OpenStudy (anonymous):

so am getting 3/2(2x+1)^-1/2-(2x+1)^-1/2 over 2x+1 is that right

OpenStudy (anonymous):

I'm not too sure if that's right maybe im just confused with the typing but I got 3(2x+1)^(1/2) - (3x+2)(2x+1)^(-1/2) all over 2x+1

OpenStudy (anonymous):

I found this really good summary of differentiation formulaes, it covers up most basic differentiation topics. Hope you find it useful for your exam coming up soon. :)

OpenStudy (anonymous):

AWESOME THANKS MAN

OpenStudy (anonymous):

how did you get rid of the 1/2 after you derived \[\sqrt{2x+1}\]

OpenStudy (anonymous):

when you differentiate root 2x+1, you use the chain rule. so its \[\sqrt{2x+1}, dy/dx= (1/2)(2)(2x+1)^{-1/2}\] so the 2 cancels out with the 1/2

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

thnxs for the help man

OpenStudy (anonymous):

no problemo ;) btw, good luck for yur exam

OpenStudy (anonymous):

thns man am gonna need it